OpenStudy (anonymous):

Solve using the addition method 2x-3y=8 3x+2y=-1

OpenStudy (dumbcow):

the goal is for one the columns, either x or y to cancel when added together look at y column -3 and 2, they will cancel if we multiply -3*2 and 2*3 now do that for the whole row 2x*2 -3y*2 = 8*2 3x*3 +2y*3 = -1*3 ___________________ 4x -6y =16 9x +6y = -3 ________________ 13x +0 = 13 x = 13/13 = 1

OpenStudy (gw2011):

2x-3y=8 Multiply this equation by 2 to eliminate y 3x+2y=-1 Multiply this equation by 3 to eliminate y 4x-6y=16 9x+6y=-3 13x=13 x=1 3(1)+2y=-1 Substituting x=1 into the second equation to find y 3+2y=-1 2y=-4 y=-2 Therefore, x=1 and y=-2

OpenStudy (anonymous):

Thank you :)

OpenStudy (gw2011):

You're welcome